# LeanQuant: Accurate and Scalable Large Language Model Quantization with Loss-error-aware Grid

### v2

@@ 3.2 Loss-Error-Aware Network Quantization
The min-max affine grid employed by previous iterative loss-error-based quantization methods do not account for the existence of inverse-diagonal outliers, which may cause significant model quality degradation. We propose loss-error-aware quantization grids, which preserve the quantized precision of weights corresponding to the inverse diagonal outliers, to produce more accurate models. We propose techniques to learn loss-error-aware grids for different quantization formats, including non-uniform and affine, and design fused GPU kernels to achieve efficient and scalable quantization.
@@ 3.2.1 Non-Uniform Loss-Error-Aware Grid
For non-uniform quantization, we perform clustering with the model parameters, weighted by their corresponding exponentiated inverse Hessian diagonals, to obtain a set of grid points that are loss-error-aware. Concretely, we learn the set of grid points $\mathcal{G}$ for $b$-bit quantization through the following objective,
@@
where $p$ is a hyperparameter that indicates the strength of preserving the quantized precision of outliers. We leverage k-means (Lloyd, 1982) to optimize the objective. After the quantization grid $\mathcal{G}$ has been established, the weights are iteratively quantized to the nearest grid point in $\mathcal{G}$.
Grid Initialization The quality of clustering results is highly dependent on initialization (Arthur et al., 2007), since Lloyd’s Algorithm (Lloyd, 1982) converges to a locally optimal solution. Hence in lower bit width regions, the initialization of grid points for clustering can greatly influence the quantized model quality. K-means++ (Arthur et al., 2007) leverages iterative sampling to select initial centers that are likely far from each other to obtain more optimal results. However, we found that it still tends to produce sub-optimal results in lower bit width regions (3 and 2-bit). We propose a lightweight and robust grid point initialization method: uniformly spaced grid initialization, which uses uniformly spaced points $\mathcal{G}_{\textrm{init}}$ over the dynamic range of the weights as initial centers, where
$\mathcal{G}_{\textrm{init}}=\Big{\{}\min(\mathbf{w})+\frac{\max(\mathbf{w})-\min(\mathbf{w})}{2^{b}-1}t\,\Big{|}\,t\in\{0,\dots,2^{b}-1\}\Big{\}}$
We perform ablative experiments to study the effectiveness of uniformly spaced grid initialization in Table 11 in the Appendix.
@@ 3.2.2 Loss-error-aware Affine Grid
We propose a different strategy for learning affine loss-error-aware grid than non-uniform grid. The challenge with learning affine grid is that the grid points are uniformly placed over an interval, so the clustering strategy does not apply. Searching for the optimal scaling factor $S$ and zero-point $Z$ over the real numbers can be done using gradient descent, but it is potentially time and memory consuming and prone to finding local minimas. Hence, we adopt the strategy of searching for the optimal $S$ and $Z$ in a constrained space $\mathbb{S}$, and formulate the problem of learning affine loss-error-aware grid as
@@
where $R=\max(\mathbf{w})-\min(\mathbf{w})$ is the dynamic range of the weights, and $T$ is the number of partitions over $R$. To obtain the $S$ and $Z$ that minimize the loss errors, we enumerate candidates for $S$ and $Z$ by shrinking the min-max range of the weights in discrete steps. The value of $T$ determines the granularity of the search, which we set to 2048.
Efficient Fused GPU Kernel for Grid Learning The learning process in Equation 7 is costly due to the enumeration of $(\frac{T}{2})^{2}$ pairs of scaling factor and zero-point. Simply looping over all enumerations is time-consuming. Hence, we design and implement a fused kernel that leverages the parallel processing power of GPUs for efficient grid searching. We designate each thread block to be responsible for processing a single group of weights, and each thread in the thread block to be responsible for enumerating all combinations of a single $t_{\mathrm{min}}$ and all $t_{\mathrm{max}}$. The minimum loss for each $t_{\mathrm{min}}$ is accumulated by each thread, and they are aggregated at the end on the thread block level to determine the optimal scaling factor and zero-point for the current weight group. This mode of processing enables the parallel computation of the optimal scaling factor and zero-point for all weight groups, accelerating the end-to-end quantization process by more than 50$\times$. An efficiency analysis of our fused kernel is presented in Section 4.3.
@@ 3.2.3 LeanQuant
Our proposed loss-error-aware quantization grid can integrate with any iterative loss-error-based quantization approaches to improve quantized model quality. A comparison between min-max affine quantization grid with loss-error-aware non-uniform and affine grid learned for a layer of Llama-3-8B (Dubey et al., 2024) is presented in Figure 1. We propose LeanQuant, which integrates loss-error-aware grid with GPTQ (Frantar et al., 2022), and present it in Algorithm 1. Additionally, for quantizing million-parameter models more accurately, we propose LeanQuant-Exact, which integrates loss-error-aware grid with OBQ (Frantar & Alistarh, 2022), and present it in Algorithm 2 in the Appendix. We further denote the underlying grid type used by LeanQuant with subscripts, such as LeanQuantaff and LeanQuantnu for affine and non-uniform grid, respectively.
The parameter $p$ balances the strength of precision preservation between inverse-diagonal outliers and non-outliers. Higher values of $p$ prioritizes the precision preservation of outliers more, while $p=0$ prioritizes the preservation of all weights equally. We found the quantized model quality to be not very sensitive to values of $p$, and use $p=4$ for all models. A sensitivity analysis for $p$ is given in Section 4.3.
@@
1:Input: weight matrix $\mathbf{W}\in\mathbb{R}^{r\times c}$, input matrix $\mathbf{X}$, bit width $b$, block size $B$, dampening factor $df$, outlier preservation strength $p$
@@
3:$\hat{\mathbf{W}}\leftarrow\mathbf{0}_{r\times c}$
4:$\mathbf{E}\leftarrow\mathbf{0}_{r\times B}$
5:$\mathbf{H}\leftarrow 2\mathbf{X}\mathbf{X}^{\top}$
6:$\mathbf{H}^{-1}\leftarrow\mathrm{Cholesky}\Big{(}\big{[}\mathbf{H}+df\cdot\mathrm{avg}\big{(}\mathrm{diag}(\mathbf{H})\big{)}\cdot\mathbf{I}\big{]}^{-1}\Big{)}$triangleright apply dampening, inversion, and Cholesky decomposition
@@
8:   $\mathcal{G}_{k}\leftarrow\operatorname*{arg\,min}\limits_{\mathcal{G}:\left|\mathcal{G}\right|=2^{b}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{nu}}(\mathbf{W}_{k,:},\mathcal{G})-\mathbf{W}_{k,:}\big{|}^{2}$ forall $k\in\{0,\dots,r-1\}$triangleright E.6
@@
10:   $S_{k},Z_{k}\leftarrow\operatorname*{arg\,min}\limits_{(S,Z)\in\mathbb{S}}\big{(}\mathrm{diag}(\mathbf{H}^{-1})^{-p}\big{)}^{\top}\big{|}\mathrm{quant}_{\textit{aff}}(\mathbf{W}_{k,:},\mathcal{S},Z)-\mathbf{W}_{k,:}\big{|}^{2}$ forall $k\in\{0,\dots,r-1\}$triangleright E.7
@@
15:         $\hat{\mathbf{W}}_{k,j}\leftarrow\mathrm{quant}_{\textit{nu}}(\mathbf{W}_{k,j},\mathcal{G}_{k})$ forall $k\in\{0,\dots,r-1\}$triangleright quantize to non-uniform grid
@@
17:         $\hat{\mathbf{W}}_{k,j}\leftarrow\mathrm{quant}_{\textit{aff}}(\mathbf{W}_{k,j},S_{k},Z_{k})$ forall $k\in\{0,\dots,r-1\}$triangleright quantize to affine grid
@@
19:      $\mathbf{E}_{:,j-1}\leftarrow\frac{\mathbf{W}_{:,j}-\hat{\mathbf{W}}_{:,j}}{\mathbf{H}^{-1}_{j,j}}$
@@
22:   $\mathbf{W}_{:,(i+B):}\leftarrow\mathbf{W}_{:,(i+B):}-\mathbf{E}\cdot\mathbf{H}^{-1}_{i:(i+B),(i+B):}$
@@ 4 Experiments
In this section, we perform extensive experiments to validate the effectiveness and scalability of our proposed LeanQuant for quantizing LLMs against competitive baselines. We first introduce the baselines, models, evaluation metrics and datasets, and hardware used for the experiments. We then describe the experimental results and findings, and analyze the efficiency and scalability of our proposal. Finally, we perform ablative experiments to validate each component of proposed approach.
Baselines We compare LeanQuantaff against competitive affine quantization approaches AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), and OmniQuant (Shao et al., 2024), and LeanQuantnu against the existing state-of-the-art non-uniform method SqueezeLLM (Kim et al., 2023). For the baselines, we use the quantized models provided by their official repository where possible, and quantize the unavailable models using their official codebase and recommended hyperparameters. More details on baseline reproduction and evaluation methods can be found in Section C of the Appendix. For all LeanQuant models, we use a small calibration set of 128 sequences of 2048 tokens from the C4 dataset (Raffel et al., 2020) for computing the Hessian $\mathbf{H}$, and set $p=4$.
Models We consider the following recent, popular LLMs for quantization: Llama 1/2/3 series models (Touvron et al., 2023a; b; Dubey et al., 2024), Mistral-7B-v0.1 (Jiang et al., 2023), Mistral-Large-Instruct-2407 (123B) (Mistral AI Team, 2024), and Llama-3.1-405B-Instruct (Dubey et al., 2024).
Evaluation Metrics and Datasets We evaluate quantized LLMs using the perplexity metric on the datasets WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), and zero-shot accuracy on the benchmarks ARC (Clark et al., 2018), LAMBADA (Paperno et al., 2016), MMLU (Hendrycks et al., 2020), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), and WinoGrande (Sakaguchi et al., 2021). We also quantize and evaluate the instruction-following Llama-3-8B-Instruct using OpenAI GPT-4o (2024-05-13) as a judge on the MT-Bench (Zheng et al., 2023), and the results are presented in Section E in the Appendix.
@@
Table 2: Zero-shot accuracy comparison of the quantized 123B Mistral-Large-Instruct-2407 model.
@@
Mistral-Large-Instruct-2407 GPTQ 4.00 84.60 63.99 74.38 80.52 76.31 77.23 89.31 85.23 78.95
LeanQuantaff 4.00 85.14 63.99 74.99 81.14 76.56 77.32 89.21 85.87 79.28
Table 3: Zero-shot accuracy comparison of the quantized Llama-3.1-405B-Instruct model.
Model Method Group Size Bits Arc-E Arc-C LAMBADA-Std PIQA Avg
Llama-3.1-405B-Instruct GPTQ 128 4.25 88.21 65.10 76.96 82.75 78.26
LeanQuantaff 128 4.25 88.34 65.70 77.86 83.03 78.73
@@ 4.1 Main Results
Accuracy and Perplexity The zero-shot accuracy of quantized models on benchmarks are presented in Table 1, as well as in Table 8 in the Appendix, and the perplexity results are shown in Table 7 in the Appendix. At the same bit width, LeanQuant achieves significantly better (lower) perplexity than GPTQ and AWQ, and performs on par with OmniQuant and SqueezeLLM. However, perplexity may not be a representative metric for evaluating the accuracy of quantized models. In terms of zero-shot accuracy on various benchmarks, LeanQuantaff mostly outperforms GPTQ and OmniQuant, and LeanQuantnu similarly performs better than SqueezeLLM in most cases. We highlight that LeanQuantaff improves the average zero-shot accuracy on 11 tasks over OmniQuant by 17.18% for 3-bit Llama-3-8B, and by 14.14% for 3-bit Mistral-7B. Compared to GPTQ, LeanQuantaff improves the average zero-shot accuracy by 18.38% for 3-bit Llama-3-8B, and by 9.16% for 3-bit Mistral-7B.
Effectiveness on Very Large LLMs We quantize the 123B Mistral-Large-Instruct-2407 and the 405B Llama-3.1 model using LeanQuantaff and GPTQ, and present their zero-shot accuracy in Table 2 and 3, respectively. OmniQuant and SqueezeLLM fail to quantize to these models due to GPU out-of-memory errors. For Llama-3.1 405B, we use a smaller set of evaluation tasks due to the high inference costs. LeanQuantaff models mostly outperforms GPTQ in zero-shot accuracy. We employ row-wise quantization for Mistral-Large and group-wise quantization (with size 128) for Llama-3.1. This showcases that our method is effective for both row-wise affine quantization and group-wise affine quantization.
@@ 4.3 Ablation Study
Q3: Is uniformly spaced grid initialization beneficial for model quality? Yes, uniformly spaced grid initialization consistently outperforms k-means++ (Arthur et al., 2007) initialization on different models in 3-bit and 2-bit regions, as shown in Table 11 in the Appendix.
Q4: Does the fused GPU kernel for LeanQuantaff accelerate quantization? Yes, our fused kernel for learning affine grids accelerate the end-to-end quantization process by more than 50$\times$, as shown in Table 5, which enables LeanQuant to be scaled to very large models.
@@
Model OmniQuant SqueezeLLM LeanQuant
Llama-3-8B 25.3 GB OOM 7.9 GB
Llama-3-70B OOM OOM 17.2 GB
Mistral-Large (123B) OOM OOM 33.0 GB
Llama-3.1-405B OOM OOM 65.4 GB

### v3

@@ 3.2 Loss-Error-Aware Network Quantization
Existing iterative loss-error-based quantization methods rely on min-max affine grids, which fail to account for outliers in the inverse Hessian diagonals. These outliers can cause significant degradation in model quality. To address this limitation, we propose loss-error-aware quantization grids that preserve the precision of weights corresponding to these outliers, thereby improving model quality. Our approach introduces techniques for learning loss-error-aware grids across various quantization formats, including non-uniform and affine. Additionally, to accelerate grid learning for large models, we developed fused GPU kernels that enable efficient and scalable quantization.
@@ 3.2.1 Non-Uniform Loss-Error-Aware Grid
For non-uniform quantization, we perform clustering on the model parameters, weighted by their corresponding exponentiated inverse Hessian diagonals, to derive a set of loss-error-aware grid points. The proposed objective aims to shape the learned grid to minimize quantization error for weights corresponding to inverse-diagonal outliers, as these outliers can disproportionately affect model quality. Concretely, we determine the set of grid points $\mathcal{G}$ for $b$-bit quantization by optimizing the following objective:
@@
Here, $p$ is a hyperparameter that balances the strength of precision preservation between inverse-diagonal outliers and non-outliers. Higher values of $p$ prioritize the precision preservation of outliers, while $p=0$ treats all weights equally. In our experiments, we set $p=4$ for all models. A sensitivity analysis for $p$ is provided in Section 4.3. To optimize this objective, we employ the k-means algorithm (kmeans), incorporating careful centroid initialization as described below. Once the quantization grid $\mathcal{G}$ is established, the weights are iteratively quantized to the nearest grid points within $\mathcal{G}$.
Grid Initialization The quality of clustering relies heavily on initialization (kmeanspp), as Lloyd’s Algorithm (kmeans) converges to a locally optimal solution. This sensitivity is especially critical in low-bit-width settings (3-bit or 2-bit), where standard methods like random or k-means++ (kmeanspp) often undersample extreme values due to the distribution of weights, which are densely concentrated near the center and sparse at the extremes.
To address this, we propose uniformly spaced grid initialization, which evenly spaces initial grid points between the minimum and maximum weights to ensure both central regions and extremes are well represented. The initial grid points for clustering, $\mathcal{G}_{\textrm{init}}$, are defined as:
$\mathcal{G}_{\textrm{init}}=\Big\{\min(\mathbf{w})+\frac{\max(\mathbf{w})-\min(\mathbf{w})}{2^{b}-1}t\,\Big|\,t\in\{0,\dots,2^{b}-1\}\Big\}$ (6)
This lightweight and robust initialization improves representation across the entire range of weights. We present an ablation study to confirm its effectiveness in Table 15 in the Appendix.
@@ 3.2.2 Loss-Error-Aware Affine Grid
The goal of learning an affine grid is to determine an optimal scaling factor $S$ and zero-point $Z$ that minimize the loss error. Unlike non-uniform grids, where clustering strategies can be applied, affine grids require the grid points to be uniformly spaced over an interval, making clustering-based approaches inapplicable. While gradient descent could be used to learn $S$ and $Z$, it is computationally intensive, memory-demanding, and susceptible to local minima.
To address this challenge, we adopt an enumerative search approach to learn the affine grid. Specifically, we enumerate candidate pairs of $S$ and $Z$ from a constrained search space $\mathbb{S}$ and select the pair that minimizes the following objective, which is similar to Equation 5:
@@
Here, $R=\max(\mathbf{w})-\min(\mathbf{w})$ is the range of the weights, $T$ is the number of partitions within $R$, and $t\in\{1,\dots,\frac{T}{2}\}$ is the number of partitions to enumerate over. By iteratively enumerating candidates of $S$ and $Z$ and evaluating their corresponding losses, we identify the optimal pair that minimizes the loss error. The parameter $T$ determines the granularity of the search; in our experiments, we set $T=2048$. To prevent overfitting, $t$ controls the amount of shrinkage of the range, which we explain in Appendix B.
Efficient Fused GPU Kernel for Grid Learning The enumerative search for $S$ and $Z$ involves evaluating $t^{2}$ candidate pairs, which can be computationally expensive if performed sequentially. To accelerate this process, we design and implement a fused GPU kernel that leverages parallel processing. Each thread block is assigned a group of weights, and individual threads within the block evaluate all combinations of a specific $t_{\mathrm{min}}$ and all possible $t_{\mathrm{max}}$. The threads compute the loss for their assigned combinations, and the results are aggregated at the block level to determine the optimal $S$ and $Z$ for the weight group. This parallelized approach enables simultaneous computation of $S$ and $Z$ across all weight groups, achieving a speedup of over $50\times$ for the end-to-end quantization process. An analysis of the kernel’s efficiency is presented in Section 4.3.
@@ 3.2.3 LeanQuant
Our proposed loss-error-aware quantization grid can be seamlessly integrated with any iterative loss-error-based quantization method to enhance the quality of quantized models. Figure 1 illustrates a comparison between the min-max affine quantization grid and loss-error-aware grids (both non-uniform and affine) applied to a layer of Llama-3-8B (llama3). We introduce LeanQuant, which combines loss-error-aware grids with GPTQ (gptq), and detail the method in Algorithm 1. Additionally, for quantizing million-parameter models more accurately, we propose LeanQuant-Exact, which integrates loss-error-aware grids with OBQ (obq), with details presented in Algorithm 2 in the Appendix. To specify the grid type used within LeanQuant, we use subscripts such as LeanQuantaff for affine and LeanQuantnu for non-uniform grids.
@@
1:Input: weight matrix $\mathbf{W}\in\mathbb{R}^{r\times c}$, input matrix $\mathbf{X}$, bit width $b$, block size $B$, dampening factor dfdf, outlier preservation strength $p$
@@
[…]
@@
[…]
@@
[…]
@@
15:   $\hat{\mathbf{W}}_{k,j}\leftarrow\mathrm{quant}_{\textit{nu}}(\mathbf{W}_{k,j},\mathcal{G}_{k})$ forall $k\in\{0,\dots,r-1\}$ $\triangleright$ quantize to non-uniform grid
@@
17:   $\hat{\mathbf{W}}_{k,j}\leftarrow\mathrm{quant}_{\textit{aff}}(\mathbf{W}_{k,j},S_{k},Z_{k})$ forall $k\in\{0,\dots,r-1\}$ $\triangleright$ quantize to affine grid
@@
19:  $\mathbf{E}_{:,j-i}\leftarrow\frac{\mathbf{W}_{:,j}-\hat{\mathbf{W}}_{:,j}}{\mathbf{H}^{-1}_{j,j}}$
@@
[…]
@@ 4 Experiments
We conduct extensive experiments to validate LeanQuant’s effectiveness and scalability in LLM quantization against competitive baselines. We first introduce the baselines, models, evaluation metrics, datasets, and hardware. Then, we present results, analyze efficiency and scalability, and conduct ablation studies to further validate our approach.
Baselines We compare LeanQuantaff against competitive affine quantization approaches AWQ (awq), GPTQ (gptq), and OmniQuant (omniquant), and LeanQuantnu against the existing state-of-the-art non-uniform method SqueezeLLM (sqllm). For the baselines, we use the quantized models provided by their official repository where possible, and quantize the unavailable models using their official codebase and recommended hyperparameters. More details on baseline reproduction and evaluation methods can be found in Section E of the Appendix. For LeanQuant models, we use a small calibration set of 128 sequences of 2048 tokens from the C4 dataset (c4) for computing the Hessian $\mathbf{H}$, and set $p=4$.
Models We consider the following recent, popular LLMs for quantization: Llama 1/2/3 series models (llama; llama2; llama3), Mistral-7B-v0.1 (mistral7b), Mistral-Large-Instruct-2407 (123B) (mistral_large), and Llama-3.1-405B-Instruct (llama3).
Evaluation Metrics and Datasets We evaluate quantized LLMs using the perplexity metric on the datasets WikiText2 (wikitext2) and C4 (c4), and zero-shot accuracy on the benchmarks ARC (arc), LAMBADA (lambada), MMLU (mmlu), HellaSwag (zellers2019hellaswag), PIQA (piqa), and WinoGrande (winogrande). We also quantize and evaluate the instruction-following Llama-3-8B-Instruct using OpenAI GPT-4o (2024-05-13) as a judge on the MT-Bench (llm_judge), and the results are presented in Section G in the Appendix.
@@
Table 2: Zero-shot accuracy of the quantized 123B Mistral-Large-Instruct-2407 model.
@@
GPTQ 4.00 84.60 63.99 74.38 80.52 76.31 77.23 89.31 85.23 78.95
Mistral-Large-Instruct-2407 LeanQuantaff 4.00 85.14 63.99 74.99 81.14 76.56 77.32 89.21 85.87 79.28
LeanQuantnu 4.05 87.67 64.59 76.63 81.51 76.50 78.00 89.35 85.68 79.99
Table 3: Zero-shot accuracy of the quantized Llama-3.1-405B-Instruct model.
Model Method Group Size Bits Arc LAMBADA MMLU Avg.
Easy Chg. Std. OpenAI STEM Human. Social Other
GPTQ 128 4.25 88.21 65.10 73.41 76.96 82.34 82.64 90.45 87.51 80.83
Llama-3.1-405B-Instruct LeanQuantaff 128 4.25 88.26 64.76 73.32 77.08 82.68 83.21 90.58 87.71 80.95
LeanQuantnu - 4.05 88.22 63.65 74.56 78.52 82.52 83.40 90.51 87.64 81.13
@@ 4.1 Main Results
Accuracy and Perplexity The zero-shot accuracy of quantized models on benchmarks are presented in Table 1, as well as in Table 10 in the Appendix, and the perplexity results are shown in Table 7 in the Appendix. At the same bit width, LeanQuant achieves significantly better (lower) perplexity than GPTQ and AWQ, and performs on par with OmniQuant and SqueezeLLM. However, perplexity may not be a representative metric for evaluating the accuracy of quantized models. In terms of zero-shot accuracy on various benchmarks, LeanQuantaff mostly outperforms GPTQ and OmniQuant, and LeanQuantnu similarly performs better than SqueezeLLM in most cases. We highlight that LeanQuantaff improves the average zero-shot accuracy on 11 tasks over OmniQuant by 17.18% for 3-bit Llama-3-8B, and by 13.38% for 3-bit Mistral-7B. Compared to GPTQ, LeanQuantaff improves the average zero-shot accuracy by 18.38% for 3-bit Llama-3-8B, and by 8.40% for 3-bit Mistral-7B.
Effectiveness on Very Large LLMs We quantize the 123B Mistral-Large-Instruct-2407 and the 405B Llama-3.1 model using LeanQuantaff, LeanQuantnu, and GPTQ, and present their zero-shot accuracy in Table 2 and 3, respectively. OmniQuant and SqueezeLLM fail to quantize to these models due to GPU out-of-memory errors. LeanQuant models mostly outperform GPTQ in zero-shot accuracy. For affine quantization, we employ row-wise quantization for Mistral-Large and group-wise quantization (with size 128) for Llama-3.1 405B. This showcases that our method is effective for both row-wise and group-wise quantization.
@@ 4.3 Ablation Study
Q3: Is uniformly spaced grid initialization beneficial for model quality? Yes, uniformly spaced grid initialization consistently outperforms k-means++ (kmeanspp) initialization on different models in 3-bit and 2-bit regions, as shown in Table 15 in the Appendix.
[…]
@@
Model OmniQuant SqueezeLLM GPTQ LeanQuant
Llama-3-8B 25.3 GB OOM 7.9 GB 7.9 GB
Llama-3-70B OOM OOM 17.1 GB 17.2 GB
Mistral-Large (123B) OOM OOM 32.8 GB 33.0 GB
Llama-3.1-405B OOM OOM OOM 65.4 GB